OpenStudy (anonymous):

describe the shape of the graph of each cubic function by determining the end behavior and number of turning points y=x^3+2x

OpenStudy (anonymous):

it is degree 3 with positive leading coefficient, so "end behaviour" is as \(x\to \infty, x^3+2x\to \infty\) and also \(x\to -\infty, x^3+2x\to -\infty\)

OpenStudy (anonymous):

this one has no turning points, since all the coefficients are positive

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

can yu try explaining it again I don't really understand it @satellite73

OpenStudy (anonymous):

do you know what i mean by "degree"?

OpenStudy (anonymous):

yes